OpenStudy (anonymous):

Verifying cot x sec^4x = cot x + 2 tan x + tan^3x ??

OpenStudy (anonymous):

I got as far as the right side =\[\frac{\cos ^{4}(x)+2\sin ^{2}(x)\cos ^{2}(x)+\sin ^{4}(x) }{ \sin(x)\cos ^{3}(x) }\] But after that I'm confused on what to do

hero (hero):

Start with the LHS

hero (hero):

\(\cot x \sec^4x = \cot x (\sec^2x)(\sec^2x)\) Right?

OpenStudy (anonymous):

no, the right hand side is all addition, not multiplication

hero (hero):

We will get there

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i get it sorry

hero (hero):

You mean you can figure it out from here?

OpenStudy (anonymous):

no im not sure on that

OpenStudy (anonymous):

i just get what you wrote

hero (hero):

\(\sec^2x = \tan^2 + 1\) correct?

OpenStudy (anonymous):

yes!

hero (hero):

So \(\cot x(\sec^2x)(\sec^2x) = \cot x(\tan^2x + 1)(\tan^2x +1)\) Do you agree?

OpenStudy (anonymous):

yeah, but where do you go from there?

OpenStudy (anonymous):

do you have to work on the other side or stay o the LHS?

hero (hero):

If we expand \((\tan^2x + 1)(\tan^2x + 1)\) what do you get from that?

hero (hero):

No, LHS only

OpenStudy (anonymous):

ok, just give me a minute to work it out

OpenStudy (anonymous):

\[cotx(\tan ^{4}x+2\tan ^{2}x+1)\] right?

hero (hero):

Looks good. Now, distribute \(\cot x\)

OpenStudy (anonymous):

should i change the tan and cot to the sin/cos and cos/sin?

hero (hero):

Nope, no need for that.

hero (hero):

All you need to know is that \(\cot x\) and \(\tan x\) are reciprocals of each other. In other words \((\cot x)(\tan x) = 1\)

hero (hero):

But if you need to see that explicitly then \(\dfrac{\cos x}{\sin x} \dot\ \dfrac{\sin x}{\cos x} = 1\)

OpenStudy (anonymous):

so if that =1, would cotx*tan^4 x= tan^3 x?

hero (hero):

Correct

OpenStudy (anonymous):

THANK YOU!

hero (hero):

You're welcome. Good luck with the rest of your work.